# KPNDepth: Depth Estimation of Lane Images under Complex Rainy Environment

### v1

@@ IV CONCLUSION
Building upon an analysis of the limitations of existing rainy weather datasets, we utilized the RCFane algorithm to create the synthetic dataset RainKITTI, capable of reflecting real rainy scenes and rain fog effects. Additionally, a new pre-trained DLKPN network for rainy image reconstruction is proposed in this paper. Integrated with traditional depth estimation models, the KPNDepth depth estimation framework is developed, employing a pixel-wise convolutional approach to restore lost depth information and address rain streak artifacts. This framework achieves more effective image reconstruction in complex rainy lane environments compared to existing methods [14][15], and outperforms the baseline model [16] in terms of accuracy in depth estimation.
Furthermore, the DLKPN framework proposed in this paper is trained on offline data, and experiments indicate its advantage in inference latency, thereby demonstrating good transferability. It offers a feasible approach for depth estimation in complex rainy environments. However, there are still some shortcomings to be addressed. For instance, our current work does not consider nighttime driving scenes, posing a research challenge on how to reconstruct the depth information of low-light objects in such scenarios.

### v2

@@ IV CONCLUSION
Our study analyzes the limitations of existing datasets and introduces the RainKITTI dataset synthesized using the RCFane algorithm to emulate real complex rainy conditions. We propose a novel pre-trained DLKPN network for rainy image reconstruction, which, when combined with conventional depth estimation models, forms the KPNDepth framework. This framework employs pixel-wise convolution to recover depth information obscured by rain streaks and enhance image reconstruction in rainy lane scenarios. It demonstrates improved performance over current methods [14][15] and surpasses the baseline model [16] in depth estimation accuracy.
Furthermore, the DLKPN framework presented in this study, trained on offline data, demonstrates superior inference latency and strong transferability, offering a viable solution for depth estimation in rainy conditions. It has been evaluated on standard rainy datasets like Rain100, yielding promising results. However, it has limitations, notably the absence of night-time rainy driving scenarios. The complex light reflections in these low-light conditions can significantly disrupt image reconstruction and depth estimation, presenting a challenge for accurately recovering depth information in such environments.
Acknowledgment
This work was partially supported by the National Science and Technology Major Project (2022ZD0117102).